# The Topology of $k$-Robust Clique Complexes in Grid-like Graphs

### v1

@@ 1 Introduction
  • Corner Attachment: When $C_{4}$ is attached along a path of length $2$, the intersection $K\cap L$ is more complex. However, we can show that $K\cap L$, while not contractible itself, is still contractible as a subcomplex of $K$. Further, $L$ is a simplicial cone $L\cong C(K\cap L)$. Then Lemma 2.17, gives $\lVert K\cup L\rVert\simeq\lVert K\vee\Sigma(K\cap L)\rVert$. Since this is not sufficient, we use the fact $k=3$ and study the homotopy type of $(K\cap L)$ further. By using a homotopy pushout–style argument, we are able to show that $\lVert K\cap L\rVert\simeq\Sigma(\bigvee_{\beta}S^{1})$, hence $\lVert\mathsf{Cliq}_{k}(H_{i+1})\rVert\simeq\lVert K\rVert\vee\Sigma^{2}(\bigvee_{\beta}S^{1})\simeq\lVert K\rVert\vee\bigvee_{\beta}S^{3}$.
@@ 3 Proofs of main results
Case 2: $G_{n}$ is attached to $H_{n-1}$ by corner gluing: We want to argue similarly as in Case 1, by describing the homotopy types of $L$ and $K\cap L$. However, the analysis is more involved.

### v2

# arXiv:2602.11365v2[math.CO] 27 Feb 2026

@@ 1 Introduction
  • Corner Attachment: When $C_{4}$ is attached along a path of length $2$, we use a slightly different decompostion: $K$ remains the same while
    $L=\{\sigma\in\mathsf{Cliq}_{k}(H_{i})\mid\sigma\cup\{x\}\in\mathsf{Cliq}_{k}(H_{i})\}.$
    One can now see that $L$ is topologically simpler - it is a simplicial cone $L\cong C(K\cap L)$. Further, we can show that $K\cap L$, while not contractible itself, is still contractible as a subcomplex of $K$. Then Lemma 2.17, gives $\lVert K\cup L\rVert\simeq\lVert K\vee\Sigma(K\cap L)\rVert$. Since this is not sufficient, we use the fact $k=3$ and study the homotopy type of $(K\cap L)$ using a homotopy pushout–style argument and we are able to show that $\lVert K\cap L\rVert\simeq\Sigma(\bigvee_{\beta}S^{1})$, hence $\lVert\mathsf{Cliq}_{k}(H_{i+1})\rVert\simeq\lVert K\rVert\vee\Sigma^{2}(\bigvee_{\beta}S^{1})\simeq\lVert K\rVert\vee\bigvee_{\beta}S^{3}$.
@@ 3 Proofs of main results
Case 2: $G_{n}$ is attached to $H_{n-1}$ by corner gluing: We want to argue similarly as in Case 1, by first decomposing $\mathsf{Cliq}_{3}(H_{n})=K\cup L$ into $K=\mathsf{Cliq}_{3}(H_{n-1})$ and $L=\{\sigma\in\mathsf{Cliq}_{3}(H_{n})\mid\sigma\cup\{x\}\in\mathsf{Cliq}_{3}(H_{n})\}$ and then describing the homotopy types of $K$,$L$ and $K\cap L$: